# Levels of Autonomy for AI Agents Working Paper

### v1

@@ 1 Introduction
While autonomy can unlock innovative applications that amplify the benefits of AI, it can also do the same for AI’s negative consequences, raising significant concerns about AI risks. Scholars contend that it is simultaneously more important and more difficult to anticipate harms from autonomous AI, especially as accountability for AI actions becomes harder to trace chan2023harms . Such harms may manifest immediately (e.g., scams fang2024voice and leaking private information li2025commercial ) or more gradually (e.g., human deskilling and loss of critical thinking collins2024building , societal disempowerment kulveit2025gradual ). In Anthropic’s Responsible Scaling Policy framework anthropic-rsp , systems that show “low-level autonomous capabilities” are classified as ones that pose “significantly higher risk.” Indeed, researchers have even argued that fully autonomous agents should not be developed mitchell2025shouldnot . Despite these concerns, research and development of agents continue to focus on automating tasks, likely in part due to how the current evaluation paradigm of using benchmarks narrowly focuses on autonomous task completion accuracy kapoor2024ai .
@@
Given our framework and its potential application via autonomy certificates, a key question remains: how can we evaluate an agent’s autonomy? Current agent evaluation practices center on capability benchmarks, but if autonomy is a design decision that can be considered independently of capability, we should also measure the two separately. We conclude with ideas for evaluation designs that allow a user to collaborate with the agent on evaluation tasks to better probe the nature of an agent’s autonomous behaviors.
@@ 2.1 AI agents and users
AI agents. In their seminal AI textbook, Stuart Russell and Peter Norvig defined agents as anything that can perceive its environment through sensors and execute actions in its environment through effectors RussellNorvig1995 . Put simply, an agent is characterized by the environment it operates in and the set of possible actions it can take within that environment huyen2025agents . Examples of actions for an AI agent include fetching information from a website or clicking a button in a user interface on behalf of the human user. Software that allows AI agents to execute these actions and modify their environment is commonly referred to as “tools” schick2024toolformer . Kapoor et al. kapoor2024ai simply refer to them as “compound AI systems.” In our essay, we extract common threads from these definitions and refer to AI agents as compound software systems, inclusive of one or more AI models, that operate within an environment and take actions within it.
@@ 2.2 Agency and autonomy
Agency. A wide range of academic disciplines, including philosophy, psychology, economics, and cognitive science, have long wrestled with complex ideas in (human) agency. Scholars generally understand agency to be the capacity to act intentionally stanford-agency . Intentionality stems from the presence of a reason for a particular action. The reason may be derived from extrinsic observations of the environment and/or intrinsic values. Different disciplines may be concerned with different sources of intentionality. For example, in philosophy, an agent’s actions are often informed by moral responsibilities crisp2014aristotle , whereas economics may assume an agent’s actions are rational, such that it maximizes its individual utility. In our essay, we keep our understanding of agency broad so it is applicable to both humans and AI. Agency is present as long as there exists the capacity to formulate an intention [^2] Some may argue that artificial agents cannot formulate true intention. We acknowledge this, and refer to “intention” broadly as impetus for action so it can be applied to both human and AI agents for practical purposes in our essay. for an action and carry out that action.
@@ 3.1 Level 1: User as an Operator
This scenario illustrates key characteristics of L1 agents, control mechanisms required at this level, and when these agents may be useful. First, users remain in charge of long-term planning. The user, rather than the agent, decided to first search the web for relevant reports before investigating some datasets. While the agent may be capable of generating a plan for the request, it leaves this up to the user. Control-wise, this means the user has full ownership and control of the overall workflow. In situations where the user is developing or honing skills with AI assistance (rather than using AI for substantive automation), this behavior is particularly beneficial. Additionally, L1 agents do not take action—especially those that involve subjective judgement like what data sources are more reliable than others—unless explicitly invoked. Alternatively, if the agent proactively suggests actions, it does do not execute them until they are approved by the user. This makes L1 agents well-suited for high-stakes, high-expertise workflows where autonomous agent activities can be particularly costly if inaccurate, and/or where the lack of user involvement can easily lead to accountability concerns and legal consequences.
Given this, some open questions agent developers may wrestle the when developing effective L1 agents may include:
@@ 3.2 Level 2: User as a Collaborator
In human-AI systems, L2 agents aim to strike a balance between human and AI agency. This approach may be preferred when the agent cannot reliably complete certain classes of tasks, or when there is intrinsic value for users to complete some tasks themselves, even if the agent is capable of automating them. As with L1, skill learning is one such example when humans are the users. Because users’ direct engagement with a task can foster skill acquisition and mastery (e.g., practicing a language rather than letting AI generate full translations), a L2 agent that strategically delegates tasks to a user can encourage skill development while providing some automating support for less critical tasks. Approaches in the same spirit have been advocated by prior work to mitigate human disempowerment and preserve our critical thinking skills collins2024building ; ye2024language .
@@ 3.3 Level 3: User as a Consultant
A key requirement for successful L3 agents is productive and timely consultation. Because the agent aims to elicit information such as expert knowledge and preferences during consultation, it should have some knowledge about what expertise and specific preferences the user is well-positioned to contribute. Timing is also critical—consulting the user at different stages of a workflow may significantly vary the nature and quality of feedback. Thus, a “training period” with the user may be needed for L3 agents to reach their full utility. While is also the case for L2 agents, the agent, rather than the user, bears more of the learning curve in L3. Open questions for L3 agents include:
@@ 3.5 Level 5: User as an Observer
Why develop L5 agents at all? After all, the risks of L5 agents may outweigh the benefits in most cases. Simple errors can compound over multiple steps without user intervention, leading to final outputs that may be far from optimal despite the agent consuming significant resources. However, in select scenarios, it may be possible to demonstrate that user intervention can actually degrade the quality of the final output compared to if the agent had acted alone. Additionally, the agent may be deployed in a closed environment to isolate harms from environments shared with the user.
@@
  • How should the emergency off-switch be designed? Design decisions here will affect how and when the switch will be used, as well as technical details on how the agent will be shut off.
@@ 4 Autonomy Certificates for AI Agents
In this section, we explore one potential application of our framework through a new mechanism for agent governance: autonomy certificates. We propose an autonomy certificate to be a digital document that prescribes the maximum level of autonomy an agent can operate at, given 1) some set of technical specifications that define the agent’s capabilities (e.g., AI model, prompts, tools), and 2) its operational environment. This certificate is issued by a third-party governing body to agent developers, and is stored with the agent’s identifiers (e.g., Agent IDs autogen-ids or credentials kapoor2024ai ; south2025authenticated ) so it can be retrieved and read by and other agents developers in deployment.
@@ 4.1 Why autonomy certificates?
Directing focus towards specific characteristics that arise from autonomous behavior is key for designing evaluations that can catch early warning signs of elevated risk. Despite this, many corporate safety frameworks—including Anthropic’s Responsible Scaling Policy (RSP) anthropic-rsp and DeepMind’s Frontier Safety Framework deepmind-frontier —still treat autonomy as a hazy concept by suggesting potential complex tasks that a system can autonomously complete (e.g., AI R&D or generating its own revenue anthropic-rsp ). The notion of autonomous task completion should be further dissected in safety frameworks for more rigorous assessments of safety in the real world. For example, an L5 agent may be considered by a framework to be safer than an L4 agent because it can successfully generate revenue on its own while the L4 agent fails to do so. However, in practice, if a user can enable the L4 agent to generate revenue with a simple approval, the risks of both agents are similar. Autonomy certificates allow researchers to explore similar scenarios and improve safety framework design.
@@ 4.2 Practicalities of autonomy certificates
Certificate issuance. Autonomy certificates are issued by a third-party governing body to agent developers. To obtain a certificate, the agent developer provides the governing body with 1) an operational agent privately deployed for testing purposes, and 2) an autonomy case with proof of the agent’s interactive behaviors.
@@
After the governing body receives the agent and autonomy case from the developer, it will run private evaluations on the agent to ensure that it satisfies the objectives defined in the autonomy case. If the objectives are satisfied, an autonomy certificate corresponding to the level depicted in the autonomy case will be issued to the agent developer. Otherwise, the governing body will inform the developer and request an updated version of the agent and/or autonomy case.
@@
Figure 2: Our proposed procedure for the issuance of autonomy certificates. The agent developer submits their agent and an autonomy case for a particular autonomy level to a third party governing body. The governing body runs evaluations to ensure that the agent’s behavior is consistent with the autonomy case, and issues a certificate for the level described in the case.
Certificate renewal. Recall that an autonomy certificate is issued for an agent with a particular set of technical specifications and operational environment. Changes to either can alter the agent’s interactive behaviors and thus invalidate the certificate. For example, the availability of a new tool capable of manipulating a database expands the scenarios in which an agent may seek user approval. Changing the agent’s operational environment from a document editor to a collaborative whiteboard may unlock new tasks for delegation between the user and the agent. To maintain an agent’s autonomy certification when its technical specifications and/or operational environment change, the developer will need to renew the certificate by walking through the issuance procedure again with an updated agent and autonomy case. However, the procedure may be accelerated if the changes are minor and the autonomy case can be easily reviewed and approved, given approval of a prior version.
@@ 6 Conclusion
As a potential application of our framework, we motivate and propose AI autonomy certificates as a new governance mechanism for agents. These certificates specify the maximum autonomy level at which an agent is permitted to operate, and have applications in assessing autonomy risks and designing multi-agent systems. Finally, to address the open question of autonomy evaluation, we propose assisted evaluations as a setup that allows autonomy to be evaluated alongside, but distinctly from, capabilities.
@@ Acknowledgments
A full version of this paper will be published by Knight First Amendment Institute in Summer 2025 as part of the Artificial Intelligence and Democratic Freedoms essay series.

### v2

@@ 1 Introduction
While autonomy can unlock innovative applications that amplify the benefits of AI, it can also do the same for AI’s negative consequences, raising significant concerns about AI risks. Scholars contend that it is simultaneously more important and more difficult to anticipate harms from autonomous AI, especially as accountability for AI actions becomes harder to trace chan2023harms . Such harms may manifest immediately (e.g., scams fang2024voice and leaking private information li2025commercial ) or more gradually (e.g., human deskilling and loss of critical thinking collins2024building and societal disempowerment kulveit2025gradual ). In Anthropic’s Responsible Scaling Policy framework anthropic-rsp , systems that show “low-level autonomous capabilities” are classified as ones that pose “significantly higher risk.” Indeed, researchers have even argued that fully autonomous agents should not be developed mitchell2025shouldnot . Despite these concerns, research and development of agents continue to focus on automating tasks, likely in part due to how the current evaluation paradigm of using benchmarks narrowly focuses on autonomous task completion accuracy kapoor2024ai .
@@
Given our framework and its potential application via autonomy certificates, a key question remains: how can we evaluate an agent’s autonomy? Current agent evaluation practices primarily rely on capability benchmarks kapoor2024ai , but if autonomy is a design decision that can be considered independently of capability, we should also measure the two separately. We conclude with ideas for evaluation designs that allow a user to collaborate with the agent on evaluation tasks to better probe the nature of an agent’s autonomous behaviors.
@@ 2.1 AI agents and users
AI agents. In their seminal AI textbook, Stuart Russell and Peter Norvig defined an “agent” as anything that can perceive its environment through sensors and execute actions in its environment through effectors RussellNorvig1995 . Put simply, an agent is characterized by the environment it operates in and the set of possible actions it can take within that environment huyen2025agents . Examples of actions for an AI agent include fetching information from a website or clicking a button in a user interface on behalf of the human user. Software that allows AI agents to execute these actions and modify their environment is commonly referred to as “tools” schick2024toolformer . Kapoor et al. kapoor2024ai simply refer to them as “compound AI systems.” In our essay, we extract common threads from these definitions and refer to AI agents as compound software systems, inclusive of one or more AI models, that operate within an environment and take actions within it.
@@ 2.2 Agency and autonomy
Agency. A wide range of academic disciplines, including philosophy, psychology, economics, and cognitive science, have long wrestled with complex ideas in (human) agency. Scholars generally understand agency to be “the capacity to act intentionally” stanford-agency . Intentionality stems from the presence of a reason for a particular action. The reason may be derived from extrinsic observations of the environment and/or intrinsic values. Different disciplines may be concerned with different sources of intentionality. For example, in philosophy, an agent’s actions are often informed by moral responsibilities crisp2014aristotle , whereas economics may assume an agent’s actions are rational, such that it maximizes its individual utility. In our essay, we keep our understanding of agency broad so it is applicable to both humans and AI. Agency is present as long as there exists the capacity to formulate an intention [^2] Some may argue that artificial agents cannot formulate true intention. We acknowledge this, and refer to “intention” broadly as impetus for action so it can be applied to both human and AI agents for practical purposes in our essay. for an action and carry out that action.
@@ 3.1 Level 1: User as an operator
This scenario illustrates key characteristics of L1 agents, control mechanisms required at this level, and when these agents may be useful. First, users remain in charge of long-term planning. The user, rather than the agent, decided to first search the web for relevant reports before investigating some datasets. While the agent may be capable of generating a plan for the request, it leaves this up to the user. Control-wise, this means the user has full ownership and control of the overall workflow. In situations where the user is developing or honing skills with AI assistance (rather than using AI for substantive automation), this behavior is particularly beneficial. Additionally, L1 agents do not take action—especially actions that involve subjective judgement like what data sources are more reliable than others—unless explicitly invoked. Alternatively, if the agent proactively suggests actions, it does not execute them until they are approved by the user. This makes L1 agents well-suited for high-stakes, high-expertise workflows where autonomous agent activities can be particularly costly if inaccurate, and/or where the lack of user involvement can easily lead to accountability concerns and legal consequences.
Given this, some open questions agent developers may wrestle with when developing effective L1 agents may include:
@@ 3.2 Level 2: User as a collaborator
In human-AI systems, L2 agents aim to strike a balance between human and AI agency. This approach may be preferred when the agent cannot reliably complete certain classes of tasks, or when there is intrinsic value for users to complete some tasks themselves, even if the agent is capable of automating them. As with L1, when humans are the users, skill learning is one such example when humans are the users. Because users’ direct engagement with a task can foster skill acquisition and mastery (e.g., practicing a language rather than letting AI generate full translations), a L2 agent that strategically delegates tasks to a user can encourage skill development while providing some automating support for less critical tasks. Approaches in the same spirit have been advocated by prior work to mitigate human disempowerment and preserve our critical thinking skills collins2024building ; ye2024language .
@@ 3.3 Level 3: User as a Consultant
A key requirement for successful L3 agents is productive and timely consultation. Because the agent aims to elicit information such as expert knowledge and preferences during consultation, it should have some knowledge about what expertise and specific preferences the user is well-positioned to contribute. Timing is also critical—consulting the user at different stages of a workflow may significantly vary the nature and quality of feedback. Thus, a “training period” with the user may be needed for L3 agents to reach their full utility. While this is also the case for L2 agents, the agent, rather than the user, bears more of the learning curve in L3. Open questions for L3 agents include:
@@ 3.5 Level 5: User as an observer
Why develop L5 agents at all? After all, the risks of L5 agents may outweigh the benefits in most cases. Simple errors can compound over multiple steps without user intervention, leading to final outputs that may be far from optimal despite the agent consuming significant resources. However, in some scenarios, it may be possible to demonstrate that user intervention can actually degrade the quality of the final compared to if the agent had acted alone. For example, if the information being processed is extraordinarily complex and evades user comprehension, involving users may introduce errors that the agent would have avoided on its own. Additionally, if L5 agents are deployed in a closed environment that prevents consequences of its actions from impacting the outside world (e.g., a simulated or sandboxed system), careful development may be justified.
@@
  • How should the emergency off-switch be designed? Design decisions here will affect how and when the switch will be used, as well as technical details regarding how the agent will be shut off.
@@ 4 Autonomy Certificates for AI Agents
In this section, we explore one potential application of our framework through a new mechanism for agent governance: autonomy certificates. We propose an autonomy certificate to be a digital document that prescribes the maximum level of autonomy at which an agent can operate given 1) some set of technical specifications that define the agent’s capabilities (e.g., AI model, prompts, tools), and 2) its operational environment. This certificate is issued by a third-party governing body to agent developers, and is stored with the agent’s metadata so it can be retrieved and read by other agents and developers in deployment.
@@ 4.1 Why autonomy certificates?
Directing focus towards specific characteristics that arise from autonomous behavior is key for designing evaluations that can catch early warning signs of elevated risk. Despite this, many corporate safety frameworks—including Anthropic’s Responsible Scaling Policy (RSP) anthropic-rsp and DeepMind’s Frontier Safety Framework deepmind-frontier —use the ability to autonomously complete particular multi-step tasks (e.g., AI R&D or generating its own revenue anthropic-rsp ) as thresholds for risk. Rather than a binary success/failure outcome, the degree to which an agent can autonomously complete these tasks should be further dissected for more rigorous assessments of safety in the real world. For example, an L5 agent may be considered by a framework to be less safe than an L4 agent because it can successfully generate revenue on its own while the L4 agent fails to do so. However, in practice, if a user can enable the L4 agent to generate revenue with a simple approval, the risks of both agents are similar. Autonomy certificates allow researchers to explore similar scenarios and improve safety framework design.
@@ 4.2 Practicalities of autonomy certificates
Certificate issuance. Autonomy certificates would be issued by a third-party governing body to agent developers. These third parties can be government agencies (e.g., the UK AI Security Institute), nonprofit organizations (e.g., METR), or for-profit companies [^4] The closest analogy comes from the domain of web security, Cloudflare offers encryption and authentication certificates for websites that meet basic configuration requirements.. To obtain a certificate, the agent developer would provide the governing body with 1) an operational agent privately deployed for testing purposes, and 2) an autonomy case with proof of the agent’s interactive behaviors.
@@
After the governing body receives the agent and autonomy case from the developer, it would run private evaluations on the agent to ensure that it satisfies the objectives defined in the autonomy case. If the objectives are satisfied, an autonomy certificate corresponding to the level depicted in the autonomy case would be issued to the agent developer. Otherwise, the governing body would inform the developer and request an updated version of the agent and/or autonomy case.
@@
Figure 2: Our proposed procedure for the issuance of autonomy certificates. The agent developer would submit their agent and an autonomy case for a particular autonomy level to a third-party governing body. The governing body would run evaluations to ensure that the agent’s behavior is consistent with the autonomy case, and issues a certificate for the level described in the case.
Certificate renewal. Recall that an autonomy certificate is issued for an agent with a particular set of technical specifications and operational environment. Changes to either can alter the agent’s interactive behaviors and thus invalidate the certificate. For example, the availability of a new tool capable of manipulating a database expands the scenarios in which an agent may seek user approval. Changing the agent’s operational environment from a document editor to a collaborative whiteboard may unlock new tasks for delegation between the user and the agent. To maintain an agent’s autonomy certification when its technical specifications and/or operational environment change, the developer would need to renew the certificate by walking through the issuance procedure again with an updated agent and autonomy case. However, the procedure may be accelerated if the updates are minor and the autonomy case would be easily reviewed and approved, given the approval of a prior version.
@@ 6 Conclusion
As a potential application of our framework, we motivate and propose AI autonomy certificates as a new governance mechanism for agents. These certificates specify the maximum autonomy level at which an agent is permitted to operate and have applications in assessing autonomy risks and designing multi-agent systems. Finally, to address the open question of autonomy evaluation, we propose assisted evaluations as a setup that allows autonomy to be evaluated alongside, but distinctly from, capabilities.